# SAM Decoding: Speculative Decoding via Suffix Automaton

### v2

@@ 3.3 Update of Suffix Automaton
After the draft is generated, we verify the draft using the large language model (LLM) and accept the correct tokens from the draft. Let the accepted tokens be denoted as $x_{\text{accept}}=(x_{t+1},x_{t+2},\ldots,x_{t+m})$. We then update the state of the suffix automaton based on these accepted tokens. For the static suffix automaton, we simply transfer the matching states according to Algorithm 1:
@@
For the dynamic suffix automaton, we first expand its state based on the accepted tokens and then transfer the matching state. Let $S_{t}$ denote the dynamic suffix automaton corresponding to the generated text $(x_{1},x_{2},\ldots,x_{t})$. The process is as follows:
$S_{t+i}$ $=\text{Expand}(S_{t+i-1},x_{t+i}),\;i\in\{1,2,...,m\},$
$s_{t+i}$ $=\text{Transfer}(S_{t+m},s_{t+i-1},x_{t+i}),\;i\in\{1,2,...,m\},$
@@ 3.4 Adaptive Draft Selection
A straightforward idea is that the length of the suffix match can indicate the quality of the draft produced by the automaton. A longer match suggests that more tokens from the draft are likely to be acceptable, whereas a shorter match implies that the draft contains less useful information. Consider that besides the retrieval-based speculative decoding approach, there exist various alternative speculative decoding techniques. Consequently, if the retrieval-based method fails to produce a satisfactory draft, we can resort to these alternatives to enhance the draft’s effectiveness.
To implement this, we concurrently employ an additional speculative decoding technique alongside the suffix automaton. During each iteration of the generation process, we adaptively select between the draft offered by the automaton and that provided by the supplementary speculative decoding method, based on the match length of the generated text within the automaton. The adaptive draft selection procedure is detailed in Algorithm 2, where $l_{\rm threshold}$ denotes a predetermined constant that determines whether to utilize the automaton or the auxiliary speculative decoding for draft generation. In our study, we incorporate two cutting-edge speculative decoding methods as auxiliary options: the model-free Token Recycling and the model-based EAGLE. Among them, Token Recycling maintains an adjacency list of the top $k$ probable next tokens for each token. During each generation step, a draft tree is built using this adjacency list via breadth-first search, with the list being continuously updated according to the latest generated tokens. Conversely, EAGLE constructs a draft tree in each generation cycle by leveraging a Transformer decoder layer that integrates the hidden states output by the LLM.
Algorithm 2 Adaptive Draft Selection
  function Drafting
     Input: static suffix automaton $S_{\rm static}$, dynamic suffix automaton $S_{\rm dyn}$, auxiliary speculative method $A$, next token $t$
     $d_{1},l_{1}=\text{Drafting}(S_{\rm static},t)$
     $d_{2},l_{2}=\text{Drafting}(S_{\rm dyn},t)$
     if $l_{1}>l_{2}+l_{\rm bias}$ then
        $d=d_{1}$
        $l=l_{1}$
     else
        $d=d_{2}$
        $l=l_{2}$
     end if
     if $l\geq l_{\rm threshold}$ then
        return $d$
     else
        return $\text{Drafting}(A,t)$
     end if
  end function
@@ 4 Experiments
Model Method Multi-turn Conversation Translation Summarization Question Answering Mathematical Reasoning Retrieval-aug. Generation #Mean Accepted Tokens Tokens/s Overall
Vicuna-7B Lookahead 1.29×1.29$\times$1.29 × 1.03×1.03$\times$1.03 × 1.22×1.$2$ 1.07×1.07$\times$1.07 × 1.41×1.41$\times$1.41 × 1.13×1.13$\times$1.13 × $1.62$ 44.37 1.19×1.19$\times$1.19 ×
REST 1.59×1.59$\times$1.59 × 1.22×1.$2$ 1.27×1.27$\times$1.27 × 1.45×1.45$\times$1.45 × 1.38×1.38$\times$1.38 × 1.36×1.36$\times$1.36 × $1.63$ 51.34 1.38×1.38$\times$1.38 ×
PLD 1.60×1.60$\times$1.60 × 0.95×0.95$\times$0.95 × 2.44×2.$4$ 1.18×1.18$\times$1.18 × 1.59×1.59$\times$1.59 × 1.72×1.72$\times$1.72 × $1.75$ 59.02 1.56×1.56$\times$1.56 ×
SAM-Decoding 1.75$\times$ 0.99$\times$ 2.54$\times$ 1.24$\times$ 1.74$\times$ 1.71$\times$ 1.89 62.26 1.65$\times$
Token Recycling 1.92$\times$ 1.61$\times$ 1.96$\times$ 1.71$\times$ 2.16$\times$ 1.68$\times$ 2.83 69.65 1.84$\times$
SAM-Decoding[T] 2.48$\times$ 1.73$\times$ 2.86$\times$ 1.98$\times$ 2.44$\times$ 2.14$\times$ 3.03 85.73 2.27$\times$
EAGLE 2.63$\times$ 1.92$\times$ 2.28$\times$ 2.11$\times$ 2.64$\times$ 1.95$\times$ 3.57 85.71 2.27$\times$
SAM-Decoding[E] 2.78$\times$ 1.88$\times$ 2.65$\times$ 2.12$\times$ 2.57$\times$ 2.10$\times$ 3.77 88.99 2.35$\times$
EAGLE2 2.87$\times$ 1.92$\times$ 2.33$\times$ 2.20$\times$ 2.88$\times$ 2.03$\times$ 4.36 90.14 2.38$\times$
SAM-Decoding[E2] 3.02$\times$ 1.89$\times$ 2.76$\times$ 2.19$\times$ 2.83$\times$ 2.23$\times$ 4.61 94.17 2.49$\times$
Model Method HumanEval MBPP HAGRID
@@
Vicuna-7B PLD 1.65 59.04 1.52×1.52$\times$1.52 × 1.42 50.62 1.27×1.27$\times$1.27 × 2.03 44.11 1.29×1.29$\times$1.29 ×
Token Recycling 2.78 75.44 1.94×1.94$\times$1.94 × 2.83 79.22 1.99×1.$9$ 2.88 66.17 1.93×1.93$\times$1.93 ×
SAM-Decoding[T] 2.94 95.08 2.45×2.45$\times$2.45 × 2.87 94.50 2.37×2.37$\times$2.37 × 3.23 87.93 2.57×2.57$\times$2.57 ×
EAGLE 4.10 103.39 2.66×2.$6$ 4.17 119.31 3.00×3.$0$ 3.44 72.14 2.11×2.$1$
SAM-Decoding[E] 4.03 110.16 2.83×2.83$\times$2.83 × 3.85 114.92 2.87×2.87$\times$2.87 × 4.05 93.17 2.72×2.72$\times$2.72 ×
EAGLE2 5.12 128.19 3.30×3.30$\times$3.30 × 5.29 135.24 3.40×3.40$\times$3.40 × 4.15 82.61 2.41×2.41$\times$2.41 ×
SAM-Decoding[E2] 4.79 127.03 3.27×3.27$\times$3.27 × 4.50 117.06 2.94×2.94$\times$2.94 × 4.75 96.60 2.81×2.81$\times$2.81 ×
Table 2: Speedup of SAM-Decoding compared to the baselines on HumanEval, MBPP, and HAGRID.
@@
Models and tasks. We conducted experiments on Vicuna-7B-v1.3 Zheng et al. (2023). We evaluated SAM-Decoding on Spec-Bench Xia et al. (2024), HumanEval Chen et al. (2021), MBPP Austin et al. (2021), and HARGID Kamalloo et al. (2023). Spec-Bench is a comprehensive benchmark designed for assessing Speculative Decoding methods across diverse scenarios. It is based on six commonly used datasets, MT-Bench Zheng et al. (2023), WMT14 DE-EN, CNN/Daily Mail Nallapati et al. (2016), Natural Question Kwiatkowski et al. (2019), GSM8K Cobbe et al. (2021), and DPR Karpukhin et al. (2020), including six aspects: Multi-turn Conversation, Translation, Summarization, Question Answering, Mathmatical Reasoning, and Retrieval-augmented Generation. In addition, MBPP, HumanEval, and HARGID are used to evaluate the speed of decoding methods in code generation tasks and contextual question answering tasks, respectively.
Baselines. We considered the following baseline methods, including the model-based method EAGLE Li et al. (2024b, a), the model-free methods Token Recycling Luo et al. (2024) and Lookahead Decoding Fu et al. (2024), and the retrieval-based methods PLD Saxena (2023) and REST He et al. (2024).
@@
Experiment setup. We conducted experiments on a server equipped with a 20-core CPU and a single NVIDIA RTX A6000 GPU (48GB). The experiments were implemented using PyTorch 2.3.0 and CUDA 12.1. For the models, we used the float16 data type and applied greedy decoding with a batch size of 1. Regarding hyperparameters, unless otherwise specified, $l_{\rm bias}$ and $l_{\rm threshold}$ were set to 5, and the size of the draft generated by the automaton was set to 40 tokens. For the auxiliary speculative decoding algorithms, we used the default configurations as described in their respective original papers. One exception was made for Token Recycling: instead of the draft tree structure used in the original paper, we employed a draft tree with a depth of 6 and containing 60 nodes. In addition, for the vicuna model, we use the vicuna template to convert prompt and output into the chat form, and for the Llama3 model, we use the template provided in the EAGLE2 repository.
@@
For SAM-Decoding, we constructed a static suffix automaton based on the Stanford-alpaca, python-code-instruction-18k, and GSK8k. To enhance our model, we incorporated two auxiliary approaches: the model-free method Token Recycling and the model-based method EAGLE. Here, SAM-Decoding[T], SAM-Decoding[E], and SAM-Decoding[E2] denote the combinations of our base model with Token Recycling, EAGLE, and EAGLE2, respectively.
Experiment Results. Experiment results on Spec-Bench when using Vicuna-7B-v1.3 are shown in Table 1. In the Spec-Bench, which consists of six tasks, multiturn conversation, summarization, and retrieval-augmented generation were identified as particularly amenable to retrieval techniques. The results indicate that integrating SAM-Decoding into these tasks led to notable speed improvements. Specifically, for Token Recycling, the speedup ratio for the three tasks rose from 1.92×1.92$\times$1.92 ×, 1.96×1.96$\times$1.96 ×, and 1.68×1.68$\times$1.68 × to 2.48×2.48$\times$2.48 ×, 2.86×2.86$\times$2.86 ×, and 2.14×2.14$\times$2.14 ×, respectively. With EAGLE, the speedup ratio increased from 2.63×2.63$\times$2.63 ×, 2.28×2.28$\times$2.28 ×, and 1.95×1.95$\times$1.95 × to 2.78×2.78$\times$2.78 ×, 2.65×2.65$\times$2.65 ×, and 2.10×2.10$\times$2.10 ×, respectively. For EAGLE2, the speedup ratios went up from 2.87×2.87$\times$2.87 ×, 2.33×2.$3$, and 2.03×2.03$\times$2.03 × to 3.02×3.02$\times$3.02 ×, 2.76×2.76$\times$2.76 ×, and 2.23×2.23$\times$2.23 ×, respectively. However, for the remaining tasks, the drafts produced by the retrieval methods did not significantly aid speculative decoding and instead introduced extra computational costs, leading to slight decreases in performance. Despite this, when SAM-Decoding was combined with these existing methods, there was an overall increase in speedup ratios: from 2.83×2.83$\times$2.83 × to 2.98×2.98$\times$2.98 × for Token Recycling, from 3.57×3.57$\times$3.57 × to 3.77×3.$7$ for EAGLE, and from 2.38×2.38$\times$2.38 × to 2.49×2.49$\times$2.49 × for EAGLE2. It is noteworthy that in the summarization task, the model-free approach (SAM-Decoding[T]) outperformed the model-based method, achieving a speedup ratio of 2.86×2.86$\times$2.86 ×. Meanwhile, these results also underscores the potential benefits of integrating SAM-Decoding with different auxiliary strategies, especially in tasks where retrieval can effectively support the decoding process.
Table 2 presents the experimental results on the HumanEval, MBPP, and HAGRID datasets when using Vicuna-7B-v1.3. For the coding datasets HumanEval and MBPP, we constructed static suffix automata based on the python-code-instructions-18k dataset. For the contextual question answering dataset HAGRID, we built a static suffix automaton using its own training set.
In the case of coding tasks, SAM-Decoding significantly accelerates the model-free method. When SAM-Decoding is combined with Token Recycling, the speedup ratios improve from 2.78×2.78$\times$2.78 × and 2.83×2.83$\times$2.83 × to 2.94×2.94$\times$2.94 × and 2.87×2.87$\times$2.87 × on the HumanEval and MBPP datasets, respectively. Notably, SAM-Decoding demonstrates a more pronounced acceleration effect on HumanEval. This is attributed to the fact that HumanEval often features longer, more semantically rich variable names and includes additional annotations in the input prompts. These conditions are favorable for retrieval-based methods. Conversely, for model-based methods like EAGLE, integrating SAM-Decoding tends to decrease the average number of accepted tokens and the generation throughput.
Turning to the contextual question answering dataset HAGRID, this task is highly conducive to retrieval methods, and SAM-Decoding effectively enhances inference speed. When SAM-Decoding is paired with Token Recycling, EAGLE, and EAGLE2, their inference speeds on the HAGRID dataset increase from 1.93×1.93$\times$1.93 ×, 2.11×2.$1$, and 2.41×2.41$\times$2.41 × to $2.57$, $2.72$, and $2.81$, respectively. It is also evident that the model-free method SAM-Decoding[T] achieves a higher speedup on this dataset compared to the state-of-the-art model-based method EAGLE2.
In addition to Vicuna-7B, we also conducted experiments on more models, and the experimental results are shown in Appendix B.
@@
Figure 3: The speedup ratio and mean accepted toknes of SAM-Decoding[T] under different draft size.
Ablation Experiments. To further understand the contributions of various components of SAM-Decoding and the influence of different hyperparameters on inference speed, we conducted a series of ablation studies. Unless otherwise stated, these experiments were carried out on Spec-Bench using the vicuna-7B-v1.3 model, with SAM-Decoding[T] serving as the decoding strategy.
Firstly, we examined the effects of $l_{\rm bias}$ and $l_{\rm threshold}$ on inference speed through a grid search. These parameters control the preference for generating text from the generating text during retrieval and the preference for using retrieval results over the draft model when creating drafts. The findings are summarized in Figure 2. As shown, both the mean accepted tokens (MAT) and the speedup ratio increase with $l_{\rm bias}$ and $l_{\rm threshold}$ up to a point. Beyond a value of 5 for both parameters, these metrics start to decline.
Additionally, we investigated how the draft size utilized by SAM-Decoding affects inference speed. Figure 3 illustrates the mean accepted tokens and the inference speed of SAM-Decoding[T] at varying draft sizes. The data indicate that as the draft size grows, the mean number of accepted tokens and inference speed initially rise. However, once the draft size surpasses 40, these metrics begin to decline, and the decrease becomes more pronounced when the draft size reaches 70.
@@
SAM-Decoding[T] 3.03 85.73 2.27×2.27$\times$2.27 ×
   w/o Static SAM 3.01 84.46 2.23×2.23$\times$2.23 ×
   w/o Dynamic SAM 2.68 76.14 2.01×2.01$\times$2.01 ×
   w/o Token Recycling 1.89 62.26 1.65×1.65$\times$1.65 ×
Finally, we investigated the impact of different modules within SAM-Decoding on inference speed. SAM-Decoding comprises three draft generation modules: the static suffix automaton, the dynamic suffix automaton, and the auxiliary speculative decoding module. We measured the inference speed of SAM-Decoding after removing each of these three modules individually. The results are presented in Table 3. From the experimental results, it is clear that each module contributes to the acceleration of the decoding process. Notably, the dynamic suffix automaton has a significantly greater impact compared to the static suffix automaton. This suggests that, in many cases, generating drafts from the dynamic context is more effective than retrieving drafts from a pre-existing text corpus. Additionally, since not all scenarios are well-suited for retrieval-based methods, the auxiliary speculative decoding module plays a crucial role in enhancing the overall performance of SAM-Decoding.
For more ablation experiment results, please refer to Appendix C.

### v3

@@ 3.3 Update of Suffix Automaton
After the draft is generated, we verify it using the large language model (LLM) and accept the correct tokens, denoted as $x_{\text{accept}}=(x_{t+1},x_{t+2},\ldots,x_{t+m})$. We then update the state of the suffix automaton based on these accepted tokens. For the static suffix automaton, we simply transfer the states according to Algorithm 1:
@@
For the dynamic suffix automaton, we first transfer the matching state based on the accepted tokens and then expand the state. Let $S_{t}$ denote the dynamic suffix automaton for the generated text $(x_{1},x_{2},\ldots,x_{t})$. The process is as follows:
$s_{t+i}$ $=\text{Transfer}(S_{t+i-1},s_{t+i-1},x_{t+i}),$
$S_{t+i}$ $=\text{Expand}(S_{t+i-1},x_{t+i}),$
$i$$\in\{1,2,...,m\},$
@@ 3.4 Adaptive Draft Selection
The retrieval-based speculative decoding methods excel at generating drafts from the corpus or the current text sequence effectively. If it fails to produce a satisfactory draft, other speculative decoding techniques can be employed to generate more diverse drafts. To combine different types of drafts, a straightforward idea is that the length of the suffix match can indicate the confidence of the draft produced by the automaton, where long matches imply that more tokens are likely to be acceptable.
To implement this, we concurrently use an auxiliary speculative decoding technique alongside the suffix automaton. During each generation step, we adaptively select the drafts offered by the automaton or the auxiliary SD method based on the match length of the generated text within the automaton. For the auxiliary SD method, we set a fixed virtual match length $l_{\rm threshold}$. In our study, we consider two auxiliary cutting-edge speculative decoding methods: the model-free Token Recycling and the model-based EAGLE-2.
Among them, Token Recycling maintains an adjacency list of the top-$k$ probable next tokens for each token and builds a draft tree using breadth-first search, and it continuously updates the list based on the latest tokens. EAGLE-2, on the other hand, leverages a Transformer decoder layer to jointly predict the last hidden states of the LLM and the next token autoregressively.
@@ 4 Experiments
Method Spec-Bench HumanEval HAGRID
@@
Lookahead* 1.63 44.37 1.20$\times$ 1.76 30.81 1.54$\times$ 1.46 23.58 1.32$\times$
REST* 1.63 51.34 1.38$\times$ 1.85 34.60 1.74$\times$ 1.53 24.91 1.39$\times$
PIA 2.08 55.45 1.47$\times$ 2.62 65.49 1.68$\times$ 2.43 66.65 1.95$\times$
PLD 1.75 59.02 1.56$\times$ 1.65 59.04 1.52$\times$ 2.03 44.11 1.29$\times$
SAM-Decoding 2.30 69.37 1.84$\times$ 2.64 88.91 2.29$\times$ 2.44 76.72 2.24$\times$
Token Recycling 2.83 69.65 1.84$\times$ 2.78 75.44 1.94$\times$ 2.88 66.17 1.93$\times$
SAM-Decoding[T] 3.03 85.73 2.27$\times$ 2.94 95.08 2.45$\times$ 3.23 87.93 2.57$\times$
EAGLE-2 4.36 90.14 2.38$\times$ 5.13 125.77 3.24$\times$ 4.15 82.61 2.41$\times$
SAM-Decoding[E2] 4.62 97.56 2.58$\times$ 4.95 130.28 3.35$\times$ 4.75 96.60 2.81$\times$
Table 1: Inference efficiency of SAM-Decoding compared to the baselines on Spec-Bench, HumanEval, and HAGRID, where * indicates that the method was compared with the baseline provided in its environment.
Figure 4: Relative speedup of SAM-Decoding compared to retrieval-based SD baselines on Spec-Bench.
Figure 5: Relative speedup of SAM-Decoding compared to SD baselines on Spec-Bench when combined with auxiliary SD methods.
@@
Models and Tasks. We conducted experiments on Vicuna-7B-v1.3 Zheng et al. (2023). We evaluated SAM-Decoding on Spec-Bench Xia et al. (2024), HumanEval Chen et al. (2021), and HARGID Kamalloo et al. (2023). Spec-Bench is a comprehensive benchmark designed for assessing Speculative Decoding methods across diverse scenarios. It is based on six commonly used datasets, MT-Bench Zheng et al. (2023), WMT14 DE-EN, CNN/Daily Mail Nallapati et al. (2016), Natural Question Kwiatkowski et al. (2019), GSM8K Cobbe et al. (2021), and DPR Karpukhin et al. (2020), including six aspects: Multi-turn Conversation (MT), Translation (Trans), Summarization (Sum), Question Answering (QA), Mathmatical Reasoning (Math), and Retrieval-augmented Generation (RAG). In addition, HumanEval, and HARGID are used to evaluate the speed of decoding methods in Code Generation task and Context Q&A task, respectively.
Baselines. We considered the following baseline methods, including the model-based method EAGLE-2 Li et al. (2024a), the model-free method Token Recycling Luo et al. (2024), and the retrieval-based methods Lookahead Decoding Fu et al. (2024), PIA Zhao et al. (2024), PLD Saxena (2023) and REST He et al. (2024).
@@
Experiment Setup. We conducted experiments on a server equipped with a 20-core CPU and a single NVIDIA RTX A6000 GPU (48GB). The experiments were implemented using PyTorch 2.3.0, Transformers 4.46.1 and CUDA 12.1. For the models, we used the float16 data type and applied greedy decoding with a batch size of 1. Regarding hyperparameters, $l_{\rm bias}$ and $l_{\rm threshold}$ were set to 5, but when there is no auxiliary method $l_{\rm bias}$ is set to 0. The size of the draft generated by the automaton was set to 40 by default, while for code datasets the size of the draft is set to 16. For the auxiliary speculative decoding methods, we used the default configurations as described in their respective original papers.
@@
For SAM-Decoding, we constructed a static suffix automaton based on the Vicuna-7B generation results on datasets Stanford-alpaca, python-code-instruction-18k, and GSK8k. To enhance our model, we incorporated two auxiliary approaches: the model-free Token Recycling and the model-based EAGLE-2. Here, SAM-Decoding[T], and SAM-Decoding[E2] denote the combinations of our base model with Token Recycling, and EAGLE-2, respectively.
Experiment Results. Experimental results on Spec-Bench, HumanEval and HAGRID when using Vicuna-7B-v1.3 are shown in Table 1. It can be seen that SAM-Decoding has higher inference speedups on all datasets compared to retrieval-based baselines, achieving speedup ratios of 1.84×1.84$\times$1.84 ×, 2.29×2.29$\times$2.29 ×, and 2.24×2.24$\times$2.24 × on each of the three datasets. Meanwhile, further speedups can be achieved by combining SAM-Decoding with other types of methods. On the Spec-Bench and HAGRID dataset, the inference speed of Token Recycling and EAGLE-2 can be further improved by combining SAM-Decoding. In Spec-Bench, the speedup ratios are improved from 1.84×1.84$\times$1.84 ×, 2.38×2.38$\times$2.38 × to 2.27×2.27$\times$2.27 ×, 2.58×2.58$\times$2.58 ×, respectively, whereas on HAGRID dataset, the speedup ratios are improved from 1.93×1.93$\times$1.93 ×, 2.41×2.41$\times$2.41 × to 2.57×2.57$\times$2.57 ×, 2.81×2.81$\times$2.81 ×. In the HumanEval dataset, the throughput of the model-based EAGLE-2 method changed slightly after integrating SAM-Decoding, due to the fact that the code generation task is less likely to copy the generated text during the generation process. Fortunately, SAM-Decoding can still speedup the model-free method Token Recycling, increasing its speedup ratios from 1.94×1.94$\times$1.94 × to 2.45×2.45$\times$2.45 ×.
In Figures 5 and 5, we further show the speedup of the different methods on each task of Spec-Bench. Compared to retrieval-based SD baselines, SAM-Decoding shows better performance across all tasks. Meanwhile, in the Spec-Bench, Multi-turn Conversation, Summarization, and Retrieval-augmented Generation were identified as particularly amenable to retrieval techniques. The results indicate that integrating SAM-Decoding into existing method led to notable speed improvements. Specifically, for Token Recycling, the speedup ratio for the three tasks raised from 1.92×1.92$\times$1.92 ×, 1.96×1.96$\times$1.96 ×, and 1.68×1.68$\times$1.68 × to 2.48×2.48$\times$2.48 ×, 2.86×2.86$\times$2.86 ×, and 2.14×2.14$\times$2.14 ×, respectively. For EAGLE-2, the speedup ratios raised from 2.87×2.87$\times$2.87 ×, 2.33×2.$3$, and 2.03×2.03$\times$2.03 × to 3.02×3.02$\times$3.02 ×, 2.76×2.76$\times$2.76 ×, and 2.23×2.23$\times$2.23 ×, respectively.
In addition to Vicuna-7B, we also conducted experiments on more models. Figure 1 shows the throughput of Vicuna-7B, Vicuna-13B and Vicuna-33B on MT-bench using SAM-Decoding and other baseline SD methods. It can be seen that SAM-Decoding outperforms retrieval-based baselines on all models. Also, SAM-Decoding can further improve the inference speed of model-free and model-based SD methods by combining them with SAM-Decoding. For more experimental results, please refer to Appendix B.
@@
Figure 7: The throughput of SAM-Decoding[T] under different draft size.
Ablation Experiments. To further understand the contributions of various components of SAM-Decoding and the influence of different hyperparameters on inference speed, we conducted a series of ablation studies.
Firstly, we examined the effects of $l_{\rm bias}$ and $l_{\rm threshold}$ on inference speed through a grid search. These parameters control the preference for generating draft from the current text over text corpus and the preference for using suffix automaton over the auxiliary SD method when creating drafts. The findings are summarized in Figure 6. We observe that both the mean accepted tokens (MAT) and the speedup ratio increase with $l_{\rm bias}$ and $l_{\rm threshold}$ before they equal 5. When the value of both parameters exceeds 5, these indicators begin to decline.
Additionally, we investigated how the draft length utilized by SAM-Decoding affects inference speed. Figure 7 illustrates the throughput of SAM-Decoding[T] at varying draft sizes. As the draft size increases, there is a positive trend in throughput until the draft size equals 40. When the draft size exceeds 40, there is an observable decline in performance metrics, which becomes more significant as the draft size reaches 70. This phenomenon can be attributed to the fact that, for draft sizes below the average acceptance length, increasing the draft size reduces the number of rounds for generation, thereby enhancing efficiency. In contrast, once the draft size surpasses this threshold, further increases do not yield additional benefits and strain GPU capacity, thus slowing inference speed.
@@
PLD 1.75 59.02 1.56$\times$
SAM-Decoding 2.30 69.37 1.84$\times$
   w/o Static SAM 1.85 61.93 1.64$\times$
   w/o Dynamic SAM 1.63 50.37 1.33$\times$
Finally, we investigated the impact of different modules within SAM-Decoding on inference speed. SAM-Decoding comprises two draft generation modules: the static suffix automaton and the dynamic suffix automaton. We measured the inference speed of SAM-Decoding after removing each of these two modules individually. The results are presented in Table 2. From the experimental results, it is clear that each module contributes to the acceleration of the decoding process. Notably, the dynamic suffix automaton has a significantly greater impact compared to the static suffix automaton. This suggests that, in many cases, generating drafts from the dynamic context is more effective than retrieving drafts from a pre-existing text corpus. For more ablation experiment results, please refer to Appendix C.